Price Quantity This is an example of a binding Price Ceiling .Economists expect that a binding Price Floor will create a Surplus in a market.TOU $90 $80 $70 $60 $50 $40 $30 $20 100 200 300 400 500 600 700 800 900 1000 Quantity a.) A price ceiling of $30 will create
Answer #1

1. Binding Price Ceiling means the government sets the minimum price for a good below the equilibrium level. It leads to a shortage in the market.

2. Binding Price Floor creates a surplus because the government sets the minimum price above the equilibrium level.

3. Shortage of 200units. i.e. (Qd-Qs = 600-400=200)

4. Shortage of 600units. i.e. (Qd-Qs = 800-200=600)

5. Surplus of 400units. i.e. (Qs-Qd = 700-300=400)